# What Guides Our Choices? Modeling Developers’ Trust and Behavioral Intentions Towards GenAI

### v1

@@ Abstract.
We developed a theoretical model to (1) identify factors that influence developers’ trust in genAI tools and (2) examine the relationship between developers’ trust, cognitive styles, and their intentions to use these tools. We surveyed software developers (N=238) at two major global tech organizations and employed Partial Least Squares-Structural Equation Modeling (PLS-SEM) to evaluate our model. Our findings reveal that genAI’s system/output quality, functional value, and goal maintenance significantly influence developers’ trust in these tools. Furthermore, developers’ trust and cognitive styles influence their intentions to use these tools. We offer practical suggestions for designing genAI tools for effective use and inclusive user experience.
@@
journalyear: 2024copyright: rightsretainedccs: Human-centered computing Empirical studies in HCI
1. Introduction
@@
Generative AI (genAI) tools (e.g., ChatGPT (OpenAI, 2024), Copilot (Microsoft, 2024)) are being increasingly used in software development (Fan et al., 2023). These tools promise enhanced productivity (Kalliamvakou, 2024) and are transforming how developers code and innovate (Peng et al., 2023). However, this push for adoption (Center, 2023) is marked with AI hype and skepticism (McKinsey, 2024), as well as interaction challenges (Fan et al., 2023; Liang et al., 2024).
@@
Trust has long been recognized as a critical design requirement of AI tools (Hoff and Bashir, 2015; Lee and See, 2004; Sellen and Horvitz, 2023). Miscalibrated levels of trust—over or under trust—can lead developers to overlook errors and risks introduced by AI (Pearce et al., 2022) or deter them from using these tools (Boubin et al., 2017). Prior research has identified various factors that foster developers’ trust in genAI tools (Wang et al., 2023; Cheng et al., 2023; Johnson et al., 2023). For instance, interaction factors such as setting appropriate expectations and validating AI suggestions (Wang et al., 2023) along with community factors like shared experiences and community support (Cheng et al., 2023) are relevant in building trust. Recently, Johnson et al. (2023) introduced the PICSE framework through a qualitative investigation with software developers, outlining key components that influence the formation and evolution of trust in software tools (see Sec. 2.1). What is missing, however, is an empirically grounded theoretical understanding of how the multitude of factors associate with developers’ trust in genAI tools. Therefore, it becomes important to answer (RQ1): What factors predict developers’ trust in genAI tools? Understanding the significance and strength of these associations is needed to inform the design and adoption of genAI tools in software development.
@@
Another important concern in industry-wide integration of AI tools is that software design can be exclusionary in different ways (Adib-Moghaddam, 2023; Beckwith and Burnett, 2004; bus, 2023), often failing to support diverse users (Eubanks, 2018). While a substantial body of work exists on modeling users’ technology acceptance (Venkatesh et al., 2003, 2012; Chau, 1996; Russo, 2024), these studies do not consider the inclusivity of the software design. One often overlooked aspect of inclusivity is supporting cognitive diversity—variations in individuals’ cognitive styles—which fosters divergence in perspectives and thoughts (see Sec. 2.2) (Sternberg and Grigorenko, 1997). Numerous studies have shown that when technology is misaligned with users’ diverse cognitive styles (Burnett et al., 2016; Beckwith and Burnett, 2004; Murphy-Hill et al., 2024), it creates additional barriers for those whose styles are unsupported, forcing them to exert additional cognitive effort (Burnett et al., 2016). Thus, it is essential to understand how developers’ cognitive styles influence their intention to adopt genAI tools, and how trust contributes to this multi-faceted decision. Therefore, we investigate (RQ2): How are developers’ trust and cognitive styles associated with their intentions to use genAI tools?
@@
We answer these research questions by establishing a theoretical model, grounded in prior literature, for trust and behavioral intentions toward genAI tools. We evaluated this model using Partial Least Squares-Structural Equation Modeling (PLS-SEM) with survey data from developers (N=238) at two major global tech organizations. Our theoretical model (Figure 1) empirically shows that genAI’s system/output quality (presentation, adherence to safe and secure practices, performance, and output quality in relation to work style/practices), functional value (educational value and practical benefits), and goal maintenance (alignment between developers’ immediate objectives and genAI’s actions) are positively associated with developers’ trust in these tools. Furthermore, developers’ trust and cognitive styles—intrinsic motivations behind using technology, computer self-efficacy within peer groups, and attitudes towards risk—are associated with their intentions to use these tools, which in turn, correlates with their reported genAI usage in work.
@@ 2.1. Trust in AI
Trust in AI is commonly defined as “the attitude that an agent will help achieve an individual’s goals in a situation characterized by uncertainty and vulnerability” (Lee and See, 2004; Liao and Sundar, 2022; Vereschak et al., 2021; Wang et al., 2023; Perrig et al., 2023). Trust is subjective and thus a psychological construct that is not directly observable (Hopkins, 1998) and should be distinguished from observable measures such as reliance (Wischnewski et al., 2023). Trust involves users attributing intent and anthropomorphism to the AI (Jacovi et al., 2021), leading to feelings of betrayal when trust is violated. Despite AI systems being inanimate, users often anthropomorphize them (Jacovi et al., 2021), thereby shifting from reliance to trust in AI systems.
@@
Unobservable psychological constructs are commonly measured through validated self-reported scales (instruments) (DeVellis and Thorpe, 2021) using questions designed to capture the construct of interest. In this paper, we measure developers’ trust in genAI tools using the validated Trust in eXplainable AI (TXAI) instrument (Perrig et al., 2023; Hoffman et al., 2023). TXAI has been derived from existing trust scales (Madsen and Gregor, 2000; Jian et al., 2000; Hoffman et al., 2023) and its psychometric quality has been validated (Perrig et al., 2023). Researchers frequently advocate using the TXAI instrument for measuring trust in AI (Perrig et al., 2023; Scharowski et al., 2024; Makridis et al., 2023).
@@
Factors affecting trust: Prior research has extensively examined factors influencing human trust in automation (Madsen and Gregor, 2000; Jian et al., 2000; McKnight et al., 2002; Merritt, 2011). However, these preliminary insights do not necessarily transfer to human-AI interactions (Wang et al., 2023) because of the nuances in how users form trust in AI tools, alongside the inherent uncertainty (Vereschak et al., 2021) and variability (Weisz et al., 2023) associated with these systems. Additionally, the context in which AI is applied (in our case, software development) influences how trust is developed and its contributing factors (Omrani et al., 2022).
@@
Relevant to our domain, Johnson et al. (2023) interviewed software engineers to outline factors that engineers consider when establishing and (re)building trust in tools through the PICSE framework: (1) Personal (internal, external, and social factors), (2) Interaction (aspects of engagement with a tool), (3) Control (over the tool), (4) System (properties of the tool), and (5) Expectations (with the tool). Since PICSE is developed for software engineering (SE), we use it to design our survey instrument to identify factors influencing developers’ trust in genAI tools. However, the PICSE framework was qualitatively developed and the psychometric quality—reliability and validity—of a survey based on it has not yet been assessed.
@@
Our work builds upon PICSE to contribute (a) a validated instrument for capturing different factors that developers consider when forming trust in genAI tools (Sec. 3.2) through a psychometric analysis of the PICSE framework and (b) assesses the significance and strength of these factors’ association with trust in genAI tools (Sec. 4).
2.2. Users’ Cognitive Styles
@@
AI can be exclusionary in different ways often failing to support all users as it should (Eubanks, 2018; Adib-Moghaddam, 2023; bus, 2023). E.g., Weisz et al. (2022) found that some, but not all, participants could produce high-quality code with AI assistance, and the differences were linked to varying participant interactions with AI.
@@
User experience in Human-AI interaction (HAI-UX) can be improved by supporting diverse cognitive styles (Anderson et al., 2022), which refer to the ways users perceive, process, and interact with information and technology, as well as their approach to problem-solving (Sternberg and Grigorenko, 1997). While no particular style is inherently better or worse, if a tool insufficiently supports (or is misaligned with) users’ cognitive styles; they pay an additional “cognitive tax” to use it, creating barriers to usability (Murphy-Hill et al., 2024).
@@
Here, we scope developers’ diverse cognitive styles to the five cognitive styles in the GenderMag inclusive design method (Burnett et al., 2016). GenderMag’s cognitive styles (facets) are users’ diverse: attitudes towards risk, computer self-efficacy within their peer group, motivations to use the technology, information processing style, and learning style for new technology. Each facet represents a spectrum. For example, risk-averse individuals (one endpoint of the ‘attitude towards risk’ spectrum) hesitate to try new technology or features, whereas risk-tolerant ones (the other end) are inclined to try unproven technology that may require additional cognitive effort or time. GenderMag’s cognitive styles are well-suited as they have been (a) repeatedly shown to align with users’ interactions with technology both in the context of SE (Murphy-Hill et al., 2024; Burnett et al., 2016; Guizani et al., 2022) and HAI interactions (Anderson et al., 2022; Hamid et al., 2024), and (b) distilled from an extensive list of applicable cognitive style types (Beckwith and Burnett, 2004; Burnett et al., 2016), intended for actionable use by practitioners. We used the validated GenderMag facet survey instrument (Hamid et al., 2023) in our study.
2.3. Behavioral Intention and Usage
@@
Behavioral intention refers to the extent to which a person has made conscious plans to undertake a specific future activity (Venkatesh et al., 2003). Technology acceptance models, such as TAM (Chau, 1996) and UTAUT (Venkatesh et al., 2003), identify behavioral intention as a key indicator of actual technology usage (Venkatesh et al., 2012). Understanding users’ behavioral intentions is useful for predicting technology adoption and guiding future design strategies (Venkatesh et al., 2003). While there is an extensive body of work modeling users’ behavioral intentions towards software tools (Venkatesh et al., 2003, 2012; Russo, 2024), these studies primarily focus on socio-technical factors driving adoption.
@@
Our work contributes to this line of research by examining the role of developers’ trust and cognitive styles in shaping their intentions to use genAI tools (Sec. 3.3 and 4), thereby extending the understanding of AI adoption dynamics in SE. We used components of the UTAUT model (Venkatesh et al., 2003) to capture developers’ behavioral intentions and usage of genAI tools.
3. Method
@@
To address our RQs, we surveyed software developers from two major global tech organizations, pseudonymized as TechNxt and TechCo. We leveraged existing theoretical frameworks and instruments to design our data collection instrument (see Sec. 2). While using existing theoretical frameworks is a first step in developing questionnaires, conducting a psychometric quality assessment is essential to ensure its subsequent reliability and validity (Furr, 2011). As there was no validated instrument to measure the constructs of the PICSE framework (Johnson et al., 2023)–our chosen trust framework–we performed its psychometric assessment (Furr, 2011) (Sec. 3.2). This assessment helped us define a theoretical model of factors developers consider when forming trust in genAI tools, which we then evaluated using Partial Least Squares-Structural Equation Modeling (PLS-SEM) to answer RQ1. To answer RQ2, we assessed the relationships between developers’ trust and cognitive styles with their intentions to use genAI tools. Next, we discuss each step.
Table 1. Measurement model constructs and instruments
@@
Trust TXAI instrument* (Perrig et al., 2023; Hoffman et al., 2023)
Factors affecting trust PICSE framework** (Johnson et al., 2023)
Users’ cognitive styles GenderMag facet survey (Hamid et al., 2023)
Behavioral intention & usage UTAUT model (Venkatesh et al., 2003, 2012)
  • *We used the 4-item TXAI scale (Hoffman et al., 2023) instead of the 6-item scale (Perrig et al., 2023) to reduce participant fatigue. **PICSE does not have a validated questionnaire in (Johnson et al., 2023).
3.1. Survey Design and Data Collection
3.1.1. Survey design
We defined the measurement model (Hair et al., 2019) based on the theoretical frameworks discussed in Sec. 2 to guide our survey design (Table 1). Four researchers with experience in survey studies and TechNxt’s research team co-designed the survey over a four-month period (Oct 2023 to Jan 2024). We adapted existing (validated) instruments in designing the survey questions (Table 1). The questions were contextualized for the target population and pragmatic decisions were made to limit the survey length. The complete questionnaire is available in the supplemental material (sup, [n. d.]).
Table 2. Respondent demographics (N=238)
@@ 3.1.1. Survey design
After the IRB-approved informed consent, participants responded to closed questions about their familiarity with genAI technology and their attitudes and intentions towards using genAI tools in work. All closed questions utilized a 5-point Likert scale ranging from 1 (“strongly disagree”) to 5 (“strongly agree”) with a neutral option. These questions also included a $6^{th}$ option (“I’m not sure”) for participants who either preferred not to or did not know how to respond to a question. This differs from being neutral–acknowledging the difference between ignorance and indifference (Grichting, 1994).
Demographic questions covered gender, continent of residence, years of software engineering (SE) experience, and primary SE responsibilities at work. We did not collect data on country of residence or specific job roles/work contexts to maintain participant anonymity, as per TechNxt and TechCo’s guidelines. An open-ended question for additional comments was included at the end of the survey.
The survey took between 7-10 minutes to complete. Attention checks were included to ensure the quality of the survey data. To reduce response bias, we randomized the order of questions within their respective blocks (each construct in Table 1). We piloted the questionnaire with collaborators at TechNxt to refine its clarity and phrasing.
3.1.2. Distribution
TechCo and TechNxt administered the online questionnaire using their internal survey tool. The survey was distributed to team leads, who were asked to cascade it to their team members. This approach was chosen over using mailing lists to ensure a broader reach (Trinkenreich et al., 2023). The survey was available for one month (Feb-Mar, 2024), and while participation was optional, it was encouraged.
3.1.3. Responses
We received a total of 343 responses: 235 from TechCo and 108 from TechNxt. We removed patterned responses (n=20), outliers ($<1$ year SE experience, n=1), and those that failed attention checks (n=29). Further, we excluded respondents who discontinued the survey without answering all the close-ended questions (n=55). We considered “I’m not sure” responses as missing data points. As in prior work (Trinkenreich et al., 2023), we did not impute data points due to the unproven efficacy of imputation methods within SEM group contexts (Sarstedt et al., 2017).
After filtration, we retained 238 valid responses (TechCo: 154, TechNxt: 84) from developers across six continents, representing a wide distribution of SE experience. Most respondents were from North America (54.2%) and Europe (23.1%), and most identified as men (78.2%), aligning with distributions reported in previous studies with software engineers (Trinkenreich et al., 2023; Russo, 2024). Table 2 summarizes the respondent demographics.
3.2. Psychometric Analysis of PICSE Framework
Psychometric quality (Lord and Novick, 2008; Raykov and Marcoulides, 2011) refers to the objectivity, reliability, and validity of an instrument. We primarily used validated instruments in designing the survey. However, since PICSE was not validated, we conducted a psychometric analysis to empirically refine its factor groupings, which were then evaluated for their association with trust (Sec. 4). Table 3 presents the factors evaluated in our survey. We performed the analysis using the JASP tool (JASP Team, 2024), adhering to established psychometric procedures (Raykov and Marcoulides, 2011; Howard, 2016; Perrig et al., 2023) as detailed next:
Table 3. PICSE framework (Johnson et al., 2023)

### v2

@@ Abstract
We developed a theoretically grounded statistical model to (1) identify factors that influence developers’ trust in genAI tools and (2) examine the relationship between developers’ trust, cognitive styles, and their intentions to use these tools in their work. We surveyed software developers (N=238) at two major global tech organizations: GitHub Inc. and Microsoft; and employed Partial Least Squares-Structural Equation Modeling (PLS-SEM) to evaluate our model. Our findings reveal that genAI’s system/output quality, functional value, and goal maintenance significantly influence developers’ trust in these tools. Furthermore, developers’ trust and cognitive styles influence their intentions to use these tools in their work. We offer practical suggestions for designing genAI tools for effective use and inclusive user experience.
Index Terms:
@@
I Introduction
@@
Generative AI (genAI) tools (e.g., ChatGPT [1], Copilot [2]) are being increasingly used in software development [3]. These tools promise enhanced productivity [4] and are transforming how developers code and innovate [5]. However, this push for adoption [6] is marked with AI hype and skepticism [7], as well as interaction challenges [3, 8].
@@
Trust has long been recognized as a critical design requirement of AI tools [9, 10, 11]. Miscalibrated levels of trust—over or under trust—can lead developers to overlook errors and risks introduced by AI [12] or deter them from using these tools [13]. Prior research has identified various factors that foster developers’ trust in genAI tools [14, 15, 16]. For instance, interaction factors such as setting appropriate expectations and validating AI suggestions [14] along with community factors like shared experiences and community support [15] are relevant in building trust. Recently, Johnson et al. [16] introduced the PICSE framework through a qualitative investigation with software developers, outlining key components that influence the formation and evolution of trust in software tools (see Sec. II-A). What is missing, however, is an empirically grounded theoretical understanding of how the multitude of factors associate with developers’ trust in genAI tools. Therefore, it becomes important to answer (RQ1): What factors predict developers’ trust in genAI tools? Understanding the significance and strength of these associations is needed to inform the design and adoption of genAI tools in software development.
@@
Another important concern in industry-wide integration of AI tools is that software design can be exclusionary in different ways [17, 18, 19], often failing to support diverse users [20]. While a substantial body of work exists on modeling users’ technology acceptance [21, 22, 23, 24], these studies do not consider the inclusivity of the software design. One often overlooked aspect of inclusivity is supporting cognitive diversity—variations in individuals’ cognitive styles—which fosters divergence in perspectives and thoughts (see Sec. II-B) [25]. Numerous studies have shown that when technology is misaligned with users’ diverse cognitive styles [26, 18, 27], it creates additional barriers for those whose styles are unsupported, forcing them to exert additional cognitive effort [26]. Thus, it is essential to understand how developers’ cognitive styles influence their intention to adopt genAI tools, and how trust contributes to this multi-faceted decision. Therefore, we investigate (RQ2): How are developers’ trust and cognitive styles associated with their intentions to use genAI tools?
We answer these research questions by establishing a theoretical model, grounded in prior literature, for trust and behavioral intentions toward genAI tools. We evaluated this model using Partial Least Squares-Structural Equation Modeling (PLS-SEM) with survey data from developers (N=238) at two global tech organizations: GitHub Inc. [28] and Microsoft[29].
@@
Our theoretical model (Figure 1) empirically shows that genAI’s system/output quality (presentation, adherence to safe and secure practices, performance, and output quality in relation to work style/practices), functional value (educational value and practical benefits), and goal maintenance (alignment between developers’ immediate objectives and genAI’s actions) are positively associated with developers’ trust in these tools. Furthermore, developers’ trust and cognitive styles—intrinsic motivations behind using technology, computer self-efficacy within peer groups, and attitudes towards risk—are associated with their intentions to use these tools, which in turn, correlates with their reported genAI usage in work.
@@ II-A Trust in AI
Trust in AI is commonly defined as “the attitude that an agent will help achieve an individual’s goals in a situation characterized by uncertainty and vulnerability” [10, 30, 31, 14, 32]. Trust is subjective and thus a psychological construct that is not directly observable [33] and should be distinguished from observable measures such as reliance [34]. Trust involves users attributing intent and anthropomorphism to the AI [35], leading to feelings of betrayal when trust is violated. Despite AI systems being inanimate, users often anthropomorphize them [35], thereby shifting from reliance to trust in AI systems.
@@
Unobservable psychological constructs are commonly measured through validated self-reported scales (instruments) [36] using questions designed to capture the construct of interest. In this paper, we measure developers’ trust in genAI tools using the validated Trust in eXplainable AI (TXAI) instrument [32, 37]. TXAI has been derived from existing trust scales [38, 39, 37] and its psychometric quality has been validated [32]. Researchers frequently advocate using the TXAI instrument for measuring trust in AI [32, 40, 41].
@@
Factors affecting trust: Prior research has extensively examined factors influencing human trust in automation [38, 39, 42, 43]. However, these preliminary insights do not necessarily transfer to human-AI interactions [14] because of the nuances in how users form trust in AI tools, alongside the inherent uncertainty [31] and variability [44] associated with these systems. Additionally, the context in which AI is applied (in our case, software development) influences how trust is developed and its contributing factors [45].
@@
Relevant to our domain, Johnson et al. [16] interviewed software engineers to outline factors that engineers consider when establishing and (re)building trust in tools through the PICSE framework: (1) Personal (internal, external, and social factors), (2) Interaction (aspects of engagement with a tool), (3) Control (over the tool), (4) System (properties of the tool), and (5) Expectations (with the tool). Since PICSE is developed for software engineering (SE), we use it to design our survey instrument to identify factors influencing developers’ trust in genAI tools. However, the PICSE framework was qualitatively developed and the psychometric quality—reliability and validity—of a survey based on it has not yet been assessed.
@@
Our work builds upon PICSE to contribute (a) a validated instrument for capturing different factors that developers consider when forming trust in genAI tools (Sec. III-B) through a psychometric analysis of the PICSE framework and (b) assesses the significance and strength of these factors’ association with trust in genAI tools (Sec. IV).
II-B Users’ Cognitive Styles
@@
AI can be exclusionary in different ways often failing to support all users as it should [20, 17, 19]. E.g., Weisz et al. [46] found that some, but not all, participants could produce high-quality code with AI assistance, and the differences were linked to varying participant interactions with AI.
@@
User experience in Human-AI interaction (HAI-UX) can be improved by supporting diverse cognitive styles [47], which refer to the ways users perceive, process, and interact with information and technology, as well as their approach to problem-solving [25]. While no particular style is inherently better or worse, if a tool insufficiently supports (or is misaligned with) users’ cognitive styles; they pay an additional “cognitive tax” to use it, creating barriers to usability [27].
@@
Here, we scope developers’ diverse cognitive styles to the five cognitive styles in the GenderMag inclusive design method [26]. GenderMag’s cognitive styles (facets) are users’ diverse: attitudes towards risk, computer self-efficacy within their peer group, motivations to use the technology, information processing style, and learning style for new technology. Each facet represents a spectrum. For example, risk-averse individuals (one endpoint of the ‘attitude towards risk’ spectrum) hesitate to try new technology or features, whereas risk-tolerant ones (the other end) are inclined to try unproven technology that may require additional cognitive effort or time. GenderMag’s cognitive styles are well-suited as they have been (a) repeatedly shown to align with users’ interactions with technology both in the context of SE [27, 26, 48] and HAI interactions [47, 49], and (b) distilled from an extensive list of applicable cognitive style types [18, 26], intended for actionable use by practitioners. We used the validated GenderMag facet survey instrument [50] in our study.
II-C Behavioral Intention and Usage
@@
Behavioral intention refers to the extent to which a person has made conscious plans to undertake a specific future activity [21]. Technology acceptance models, such as TAM [23] and UTAUT [21], identify behavioral intention as a key indicator of actual technology usage [22]. Understanding users’ behavioral intentions is useful for predicting technology adoption and guiding future design strategies [21]. While there is an extensive body of work modeling users’ behavioral intentions towards software tools [21, 22, 24], these studies primarily focus on socio-technical factors driving adoption.
@@
Our work contributes to this line of research by examining the role of developers’ trust and cognitive styles in shaping their intentions to use genAI tools (Sec. III-C and IV), thereby extending the understanding of AI adoption dynamics in SE. We used components of the UTAUT model [21] to capture developers’ behavioral intentions and usage of genAI tools.
III Method
@@
To address our RQs, we surveyed software developers from two major global tech organizations, GitHub Inc. and Microsoft. We leveraged existing theoretical frameworks and instruments to design our data collection instrument (see Sec. II). While using existing theoretical frameworks is a first step in developing questionnaires, conducting a psychometric quality assessment is essential to ensure its subsequent reliability and validity [51]. As there was no validated instrument to measure the constructs of the PICSE framework [16]–our chosen trust framework–we performed its psychometric assessment [51] (Sec. III-B). This assessment helped us define a theoretical model of factors developers consider when forming trust in genAI tools, which we then evaluated using Partial Least Squares-Structural Equation Modeling (PLS-SEM) to answer RQ1. To answer RQ2, we assessed the relationships between developers’ trust and cognitive styles with their intentions to use genAI tools. Next, we discuss each step.
III-A Survey Design and Data Collection
III-A1 Survey design
We defined the measurement model [52] based on the theoretical frameworks discussed in Sec. II to guide our survey design (Table I). Four researchers with experience in survey studies and GitHub’s research team co-designed the survey over a four-month period (Oct 2023 to Jan 2024). We adapted existing (validated) instruments in designing the survey questions (Table I). The questions were contextualized for the target population and pragmatic decisions were made to limit the survey length. The complete questionnaire is available in the supplemental material [53].
TABLE I: Measurement Model Constructs and Instruments
@@ III-A1 Survey design
Trust TXAI instrument* [32, 37]
Factors affecting trust PICSE framework** [16]
Users’ cognitive styles GenderMag facet survey [50]
Behavioral intention & usage UTAUT model [21, 22]
  • *We used the 4-item TXAI scale [37] instead of the 6-item scale [32] to reduce participant fatigue. **PICSE does not have a validated questionnaire in [16].
After the IRB-approved informed consent, participants responded to closed questions about their familiarity with genAI technology and their attitudes and intentions towards using genAI tools in work. All closed questions utilized a 5-point Likert scale ranging from 1 (“strongly disagree”) to 5 (“strongly agree”) with a neutral option. These questions also included a $6^{th}$ option (“I’m not sure”) for participants who either preferred not to or did not know how to respond to a question. This differs from being neutral–acknowledging the difference between ignorance and indifference [54].
Demographic questions covered gender, continent of residence, years of software engineering (SE) experience, and primary SE responsibilities at work. We did not collect data on country of residence or specific job roles/work contexts to maintain participant anonymity, as per GitHub and Microsoft’s guidelines. An open-ended question for additional comments was included at the end of the survey.
The survey took between 7-10 minutes to complete. Attention checks were included to ensure the quality of the survey data. To reduce response bias, we randomized the order of questions within their respective blocks (each construct in Table I). We piloted the questionnaire with collaborators at GitHub to refine its clarity and phrasing.
TABLE II: Respondent Demographics (n=238)
@@
III-A2 Distribution
GitHub and Microsoft administered the online questionnaire using their internal survey tools. The survey was distributed to team leads, who were asked to cascade it to their team members. This approach was chosen over using mailing lists to ensure a broader reach [55]. The survey was available for one month (Feb-Mar, 2024), and while participation was optional, it was encouraged.
III-A3 Responses
We received a total of 343 responses: 235 from Microsoft and 108 from GitHub. We removed patterned responses (n=20), outliers ($<1$ year SE experience, n=1), and those that failed attention checks (n=29). Further, we excluded respondents who discontinued the survey without answering all the close-ended questions (n=55). We considered “I’m not sure” responses as missing data points. As in prior work [55], we did not impute data points due to the unproven efficacy of imputation methods within SEM group contexts [56].
After filtration, we retained 238 valid responses (Microsoft: 154, GitHub: 84) from developers across six continents, representing a wide distribution of SE experience. Most respondents were from North America (54.2%) and Europe (23.1%), and most identified as men (78.2%), aligning with distributions reported in previous studies with software engineers [55, 24]. Table II summarizes the respondent demographics.
III-B Psychometric Analysis of PICSE Framework
Psychometric quality [57, 58] refers to the objectivity, reliability, and validity of an instrument. We primarily used validated instruments in designing the survey. However, since PICSE was not validated, we conducted a psychometric analysis to empirically refine its factor groupings, which were then evaluated for their association with trust (Sec. IV). Table III presents the factors evaluated in our survey. We performed the analysis using the JASP tool [59], adhering to established psychometric procedures [58, 60, 32] as detailed below:
[…]
TABLE III: PICSE framework [16]